MEGHÍVÓ

2004. április 30-án, pénteken, a Nagyteremben 15 óra 30 perckor kezdődő Összintézeti Szemináriumunk programja:

JEAN-CHRISTOPHE YOCCOZ (Collège de France)

INTERVAL EXCHANGE MAPS

című előadása.

Interval exchange maps are natural generalizations of rotations on a circle. A continuous fraction algorithm, developped by Rauzy, Veech and Zorich, allows to characterize a diophantine property under which the natural difference equation may be solved. This is joint work with S. Marmi and P. Moussa.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!

2004. április 30-án, pénteken, a Nagyteremben 15 óra 30 perckor kezdődő Összintézeti Szemináriumunk programja:

JEAN-CHRISTOPHE YOCCOZ (Collège de France)

INTERVAL EXCHANGE MAPS

című előadása.

Interval exchange maps are natural generalizations of rotations on a circle. A continuous fraction algorithm, developped by Rauzy, Veech and Zorich, allows to characterize a diophantine property under which the natural difference equation may be solved. This is joint work with S. Marmi and P. Moussa.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!